OpenStudy (anonymous):

what is the range of the following function y=1/(x+3)+2

OpenStudy (anonymous):

do you know the range of y=1/x ?

OpenStudy (anonymous):

would it be 0?

OpenStudy (anonymous):

the graph of your function, \(\ y= \frac{1}{x+3}+2 \) , is a translation of y=1/x three units left, 2 units up. so if you know the range of y=1/x, the range is just a matter of putting the "2" in there.

OpenStudy (anonymous):

yes, all real numbers except 0... if that's what you meant....

OpenStudy (anonymous):

oh okay so that the answer ? or do i still have to put the "2" in there?

OpenStudy (anonymous):

since your function is 2 units "higher", your range would be : all real numbers EXCEPT 2

OpenStudy (anonymous):

ohhh okay

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

yw...:)

OpenStudy (anonymous):

The range of a function are all the numbers that you get out if you plug in all values of x. If you look at the graph of this function, you can see that all real numbers on the y axis have some number x that corresponds to them. The one exception is y=2. Is there any value of x that you can plug into f(x) and still get y=2?